### v2

# Proper Body Landmark Subset Enables More Accurate and 5X Faster Recognition of Isolated Signs in LIBRAS

Abstract
[…]
1 Introduction
Figure 1: Pipeline of the proposed landmark-based approach for isolated sign recognition in LIBRAS. Video frames are processed with MediaPipe to extract landmarks, which are selected (subset selection), interpolated (imputation), encoded as 2-D skeleton images, and classified by a CNN to predict the sign label.
Artificial intelligence has enabled significant advancements in accessible technologies across various domains (Hossain et al., , 2023; Pereira et al., , 2025; Pantera et al., , 2025). In particular, there is a growing demand for solutions that support individuals with partial or total hearing loss, a population estimated at $1.57$ billion in 2019, and projected to reach $2.45$ billion worldwide by 2050 (Guo et al., , 2024; Nieman et al., , 2024).
In this context, several works have addressed sentence-level (continuous) (Wang et al., , 2025; Hu et al., , 2023) and word-level (isolated) (Alves et al., , 2024; Kumari and Anand, , 2024) recognition of Sign Language (SL) from video sequences. In the latter, which is the focus of our investigation, the objective is to identify the sign (“word”) performed by an individual (signer) recorded in front of a camera, assuming that a single sign is performed. This application is particularly helpful for self-paced SL training platforms (Starner et al., , 2024), visual keyword search for SL (Tamer and Saraçlar, , 2020), dictionary lookup by example (Bohacek and Hassan, , 2023), as well as query-by-example sign spotting to locate occurrences in continuous videos (Varol et al., , 2022; Vázquez Enríquez et al., , 2022), and cross-modal retrieval to search sign videos from free-form text (Duarte et al., , 2022).
Most works addressing Isolated Sign Language Recognition (ISLR) from video sequences leverage deep neural networks, such as 2- and 3-D Convolutional Neural Networks (CNNs) (Alves et al., , 2024; De Castro et al., , 2023), Gated Recurrent Units (GRUs) (Shen et al., , 2024), Long Short-term Memory (LSTM) networks (Aly and Aly, , 2020; Rastgoo et al., , 2021), and Transformers (Ertürk et al., , 2025; Boháček and Hrúz, , 2022; Pu et al., , 2024). Promising results have been achieved by combining multiple modalities. For instance, Rastgoo et al., (2021) employed a 2-D CNN to extract four types of spatial features: global frame features, local hand features, optical flow, and hand-pose features. These features were fused with depth maps generated by Microsoft Kinect, and subsequently fed into an LSTM for temporal processing.
Multimodality was also explored by De Castro et al., (2023), who employed 3-D CNNs across different scenarios, ranging from the use of raw RGB data to the integration of optical flow, synthetically generated depth maps, segmented hands and faces, and distance/velocity maps derived from body landmarks (joints and other keypoints). Their best results were achieved with a complex multi-stream architecture in which each branch processed: (i) joint distance and velocity maps; (ii) segmented hands; (iii) segmented faces; and (iv) the original RGB frames augmented with a depth map. Despite its effectiveness, the approach introduces significant complexity and computational overhead, as landmark extraction depends on OpenPose (Cao et al., , 2021), a framework that delivers good accuracy but at a high computational cost.
To address the complexity of the multi-stream design, Alves et al., (2024) proposed a single-stream pipeline where body landmarks – covering the trunk, hands, and face – are extracted from RGB frames and transformed into a 2-D skeleton-image representation (Memmesheimer et al., , 2022), which is then classified by a 2-D CNN. Their approach outperformed De Castro et al., (2023), achieving state-of-the-art recognition accuracy on MINDS-Libras (Rezende et al., , 2021) and LIBRAS-UFOP (Cerna et al., , 2021), the most widely used LIBRAS benchmarks in the context of ISLR. However, their method still relies on OpenPose for landmark extraction, incurring considerable computational cost and posing a barrier to real-time applications.
This limitation has motivated the exploration of lightweight frameworks for landmark extraction, such as MediaPipe (Lugaresi et al., , 2019), enabling greater efficiency and scalability. Ertürk et al., (2025) used MediaPipe to extract landmarks, which are processed by a Transformer network (Vaswani et al., , 2017), achieving over 90% accuracy in Turkish SL. Similarly, Laines et al., (2023) convert MediaPipe holistic landmarks (hands, face, and upper body) into Tree-Structure Skeleton Images and report state-of-the-art results among skeleton-based models on WLASL, with competitive performance on AUTSL. Luna-Jiménez et al., (2023) employed a Transformer network over MediaPipe landmarks (hands + pose). They showed that using MediaPipe improves WLASL100 accuracy by 13.18 percentage points compared to non-MediaPipe features (Luna-Jiménez et al., , 2023).
The time efficiency enabled by MediaPipe has motivated us to investigate its impact on the ISLR framework proposed by Alves et al., (2024). However, preliminary exploration showed that using all landmark points provided by MediaPipe dramatically reduced recognition accuracy. Based on this observation, we conducted this investigation to verify whether landmark subsets used in the literature, as well as in winning solutions publicly available from SLR challenges, enable the use of MediaPipe for MINDS-Libras and LIBRAS-UFOP. The main findings in this paper are:
  • Selecting an appropriate subset of landmarks substantially improves recognition performance, achieving higher accuracy while reducing input dimensionality.
  • Applying a simple spline-based imputation for missing landmarks – common in MediaPipe detections – further enhances accuracy.
  • The proposed method, using the best-performing landmark strategy, outperforms existing approaches in the literature.
[…]
The remainder of this paper is organized as follows. Section 2 presents the landmark-based framework for recognition of isolated signs. Section 3 describes the experimental methodology, while the results are presented and analyzed in Section 4. Finally, Section 5 discusses the concluding remarks and directions for future work.
2 Landmark-based ISLR Framework
The landmark-based ISLR pipeline is illustrated in Figure 1. The input sample is a video sequence associated with a single sign. Each RGB frame of the video is processed individually to detect body landmarks. This procedure relies on the MediaPipe Holistic model (Lugaresi et al., , 2019), which combines pose, face, and hand markers into a unified full-body estimation. The detected landmark points are organized into a CSV file, where the columns represent the xx/$y$ normalized coordinates of each specific point, and the rows represent the temporal sequence.
A subset of the detected landmarks (i.e., columns of the CSV file) is selected according to a predefined strategy. Due to noise in the landmark detection process, several missing points can be observed in the CSV file. To address this issue, we applied spline-based interpolation to impute the missing points. This approach shortens detection gaps, yielding temporally continuous trajectories for each point. The processed landmarks are subsequently converted into a 2-D skeleton image representation (Memmesheimer et al., , 2022), which compactly encodes the spatial structure and temporal dynamics of the landmarks. Finally, the image is resized to $224\times 224$ and fed into a CNN (ResNet-18), which outputs the sign label based on the highest probability.
The following subsections detail each stage of the pipeline: landmark extraction, landmark subset selection, imputation, image encoding, and the training process used to deploy the classification model.
2.1 Landmarks Extraction
Figure 2: Landmarks subset selection: the columns show the set of landmarks utilized in each each strategy and the respective count across the body parts.
The MediaPipe framework has been extensively adopted in real-time human perception tasks (Uddin et al., , 2025; Sultana et al., , 2024; Bagga and Yang, , 2024), due to its ability to capture manual articulators (hands), upper-body kinematics, and orofacial cues. Our ISLR pipeline relies on the Holistic model, which outputs full-body landmarks, including (by default) 543 points (Figure 2 (baseline)): 468 for the face, 33 for the pose (estimating body posture per frame across the entire body), and 21 per hand (42 in total). Note that the pose landmarks include a few coarse points around the head (e.g., nose and eyes), but the detailed facial landmarks are provided as a separate component. MediaPipe also provides depth information ($z$-coordinate) and visibility scores. Our approach, however, operates strictly in 2-D and therefore discards the additional information provided by the MediaPipe framework.
Settings
The minimum detection confidence and minimum tracking thresholds in MediaPipe were set to 0.4. This value proved effective in preliminary tests for stabilizing landmark extraction. The former triggers the initial detection of the articulators (e.g., hands, face, and pose) in each frame, while the latter controls temporal tracking to maintain consistent trajectories across frames.
2.2 Landmarks Subset Selection
As mentioned, 468 out of the 543 landmarks provided by the Holistic model correspond to the face alone, potentially introducing redundancy (non-linguistic variation) and noise. Based on the premise that appropriate landmark subsets are key to improving recognition accuracy, five strategies for landmark selection are considered in this study: the full set of points (baseline) and four subsets focusing on the most relevant articulators for SLR. The landmark (sub)sets are described in the following.
Baseline (All)
This strategy (Figure 2, col. 1) uses all 543 landmarks provided by MediaPipe Holistic, serving as a reference for comparison against the other strategies evaluated in this work. By retaining the full configuration, it maximizes spatial and articulatory information, preserving facial micro-configurations (mouth, eyebrows, and gaze), upper-limb kinematics, and manual gestures. The trade-off is a substantially larger input data and increased susceptibility to noise.
Laines et al., (2023)
This strategy (Figure 2, col. 2) employs 68 landmarks distributed across the face, pose, and hands, prioritizing regions functionally relevant for signing (lips/expressivity, shoulder girdle and upper-limb joints, plus key points on both hands). As noted by the authors, the goal is to reduce dimensionality while preserving the articulatory mechanisms most critical for ISLR: manual configurations, arm/forearm orientation and perspective, and facial cues related to the mouth and expressiveness. Compared to the baseline (All), this approach is expected to provide greater robustness to noise and lower computational cost, with potential performance gains when signs are strongly driven by the hands and salient facial/head regions.
Arcanjo et al., (2024)
This configuration (Figure 2, col. 3) retains pose and hands while entirely excluding the dense points of the face. It uses 33 pose points and 42 hand points, totaling 75 landmarks. As seen in Figure 2, there are coarse points on the face provided by the pose component of MediaPipe. Removing the dense points from the face results in a substantial reduction in dimensionality, simplifying the hypothesis space and aiding regularization.
ASL Signs Challenge 1st-place (ASL-1st)
This strategy (Figure 2, col. 4) adopts the subset from the winning solution (Sohn, , 2023) of the Google ASL Signs Challenge on Kaggle, which focuses on isolated signs. A total of 118 landmarks were selected from the face and hands, excluding body pose. The emphasis is on manual configurations and the orofacial region (lips/jaw), which are distinctive in many lexicons, while global pose is discarded to reduce dimensionality and dependence on posture. The underlying observation is that, for many isolated signs, the differential cues lie in hand shape, contact/relative position to the face, and lip articulation rather than in trunk alignment.
ASL Signs Challenge 2nd-place (ASL-2nd)
Also derived from the Google ASL Signs Challenge on Kaggle, this configuration (Figure 2, col. 5) replicates the subset from the runner-up solution (Toporov et al., , 2023), comprising 80 keypoints concentrated on the lips, hands, and body pose. Unlike ASL-1st, it preserves a core of pose landmarks to retain postural context (shoulder, elbow, and wrist angles) while still prioritizing manual articulation and the labial region. The aim is to balance fine-grained hand discrimination with sufficient global information to disambiguate signs with distinct trajectories or ranges.
2.3 Spline-Based Landmark Imputation
The imputation procedure addresses detection gaps resulting from MediaPipe failures. To this end, we apply piecewise spline interpolation, treating each landmark-coordinate combination across time as an independent function to be reconstructed. Missing values are estimated from their short temporal neighborhoods (window size of 5), leveraging information from both past and future frames. Specifically, we employ a cubic spline for robust reconstruction when at least four points are available in the series; otherwise, the procedure defaults to linear interpolation. The short-window constraint preserves kinematic plausibility and prevents extrapolation beyond the observed range. This step stabilizes the spatiotemporal trajectories prior to the encoding stage, mitigating discontinuities that could otherwise degrade classifier training.
2.4 Image Encoding
Landmarks are encoded as images using SkeletonDML, an algorithm proposed by Memmesheimer et al., (2022), originally applied to represent skeletal joints in few-shot action recognition. To the best of our knowledge, Alves et al., (2024) were the first to apply Skeleton-DML to ISLR, obtaining promising results on LIBRAS datasets. The underlying idea of such representation applies broadly to space–time coordinates, such as general body landmarks. Adapting it to the 2-D case is straightforward by disregarding depth ($z$-coordinate), which simply reduces the encoded data volume.
Given $L$ landmarks and $T$ frames, the 2-D version of the Skeleton-DML encoding procedure consists of creating two $L\times T$ matrices (tensors), one for each of the $x$ and $y$ coordinates. These tensors are reshaped so that three consecutive column entries are organized as channels, forming a 3-channel image. The reshaped tensors are then horizontally concatenated, forming a single image. More details are provided in the original work (Memmesheimer et al., , 2022). For a more formal description, the reader is referred to Alves et al., (2024).
2.5 Training Protocol
The training protocol follows the procedure utilized by Alves et al., (2024), with a single modification, that we train for 30 epochs instead of the original 20. At every epoch, we perform on-the-fly data augmentation directly on the landmark sequences prior to their encoding into 2-D images, consisting of rotation, zoom, translation, and horizontal flip to increase data diversity without changing the number of samples.
The classifier model is a ResNet-18 initialized with ImageNet weights. Its final convolutional output is first passed through a global average pooling layer, producing a 512-dimensional feature vector. This vector is then fed into two fully connected layers: a 128-unit layer with ReLU activation, followed by an output layer matching the number of classes. The network input is $224\times 224$, while landmark-encoded images have size $126\times(2T/3)$, where $T$ is the number of frames of that particular input video. For the evaluated datasets, $2T/3<224$, so resizing to $224\times 224$ results in upsampling without information loss.
Optimization leverages Adam to minimize cross-entropy, with batch size $64$, learning rate $10^{-4}$, and 30 epochs. Early stopping is triggered if the validation loss fails to decrease for five consecutive epochs. To prevent overfitting, we apply (i) batch normalization after the 512-dimensional layer, (ii) dropout (with 50% probability) after the 128-unit layer, and (iii) L2 regularization (weight decay of $10^{-4}$) on all trainable parameters. The model selected for evaluation is the checkpoint achieving the highest validation accuracy.
3 Experimental Methodology
This section outlines the experimental framework used for performance assessment, including the Brazilian Sign Language (LIBRAS) datasets, evaluation metrics, experiments, and the computational setup.
3.1 Datasets
Two public popular ISLR LIBRAS datasets were considered in this investigation: MINDS-Libras (Rezende et al., , 2021) and LIBRAS-UFOP (Cerna et al., , 2021). In particular, these collections enable direct comparison with Alves et al., (2024), whose methodology motivated this study.
3.1.1 MINDS-Libras
This collection comprises 1,155 video recordings – each representing a sign – covering 20 signs selected according to phonological parameters (handshape, location, orientation, movement, and non-manual markers). The vocabulary includes “To happen”, “Student”, “Yellow”, “America”, “To enjoy”, “Candy”, “Bank”, “Bathroom”, “Noise”, “Five”, “To know”, “Mirror”, “Corner”, “Son”, “Apple”, “Fear”, “Bad”, “Frog”, “Vaccine”, and “Will”. The executions were performed by 12 signers varying in sex, age, and fluency in LIBRAS. Although clothing colors were not strictly standardized, most signers wore dark (black) attire, minimizing interference with the background.
Data collection was conducted in a controlled environment with a fixed chroma key background, and the sensors were positioned to capture the movement of the upper limbs. Two acquisition modalities were employed: an RGB camera (Canon EOS Rebel T5i, resolution 1920 $\times$ 1080) and an RGB-D sensor (Kinect v2, depth resolution 640 $\times$ 480), with only the RGB modality leveraged by our method. The typical recording distances were approx. 1.60m for the RGB-D sensor and approx. 1.92m for the RGB camera.
3.1.2 LIBRAS-UFOP
This is a multimodal dataset created with a Microsoft Kinect v1 device that captures synchronized recordings in three modalities: RGB, depth (RGB-D), and 3D skeleton data. Similar to MINDS-Libras, only the RGB modality is explored in our experiments. The dataset comprises 56 signs, carefully selected, grouped, and organized into four categories to emphasize fine-grained phonological distinctions.
The recordings were performed by five signers (three female and two male). To increase variability, data collection took place in two different scenarios, under diverse illumination conditions, and without strict standardization of the distance between the participant and the sensor. In addition, subjects executed the signs at different speeds and considered both one-handed and two-handed executions, depending on the lexical nature of the sign. Data capturing was conducted at 30 fps, with resolution of 640 $\times$, 480 and a total of 3,040 sequences were obtained. A particular feature of this dataset is that each video comprises between 8 and 16 repetitions of the same sign performed by a given signer. Since the recognition pipeline requires each sign to be individualized, we use the same time cut points provided by Alves et al., (2024).
3.2 Recognition Performance Metrics
Recognition performance was measured in terms of accuracy, as well as macro-averaged precision, recall, and F1-score. These are traditional metrics in the ISLR domain (De Castro et al., , 2023; Alves et al., , 2024) derived from the counts of the confusion matrix in a multiclass context: TP (true positives), instances that belong to the target class and are correctly predicted as such; TN (true negatives), instances correctly predicted as not belonging to the target class; FP (false positives), instances that do not belong to the target class but are incorrectly predicted as belonging; and FN (false negatives), instances that belong to the target class but are incorrectly predicted as not belonging.
Accuracy quantifies the proportion of correct predictions over the total number of instances and serves as a global indicator of performance (Sokolova and Lapalme, , 2009; Powers, , 2011). Precision measures the reliability of positive predictions, while recall (sensitivity) measures the coverage of actual positives. To balance these aspects, the F1-score –the harmonic mean of precision and recall – often provides a more informative summary than accuracy in the presence of class imbalance. The use of precision, recall, and F1-score enables a more comprehensive evaluation and allows direct comparison with state-of-the-art methods, although no significant imbalance was observed in either evaluated dataset (see Figure 3).
Figure 3: Distribution of video sequence length w.r.t. signs for the evaluation datasets.
3.3 Experiments
The conducted experiments address one main research question and three additional ones:
  • RQ1 (Main Question): What is the impact of landmark selection on recognition performance?
  • RQ2: What is the impact of the spline-based imputation method employed in our approach?
  • RQ3: How does our method, configured with the best-performing strategy, compare with existing approaches in the literature (Alves et al., , 2024; De Castro et al., , 2023; Cerna et al., , 2021; Passos et al., , 2021)?
  • RQ4: What speed-up does our MediaPipe-based solution provide compared to the OpenPose-based pipeline by Alves et al., (2024)?
3.3.1 Main Experiment (RQ1 + RQ2)
For each video, the full set of 543 landmarks was extracted using the MediaPipe Holistic pipeline. We retained the original raw data stream, containing the spatiotemporal trajectories of all points, as well as a processed stream obtained after the spline-based imputation (Section 2.3) to mitigate tracking errors and detection gaps. Note that both streams comprise the full set of detected landmarks, serving to establish a comparison baseline. From each stream, we derived four additional ones by applying the selection strategies introduced in Section 2.2 (Laines, Arcanjo, ASL-1st, and ASL-2nd), resulting, therefore, in 10 streams to be tested. These streams were encoded into $224\times 224\times 3$ skeleton images, following the method presented in Section 2.4.
Finally, each encoded image set was subjected to train-test sessions (Section 2.5). Following best practices (Alves et al., , 2024; De Castro et al., , 2023), we adopted a nested Leave-One-Person-Out (LOPO) cross-validation protocol, where data from one signer is used for testing while the remaining n1n-1 signers’ data are used for training and validation. In the nested approach, for a given signer in the test set, each of the n1n-1 remaining signers is detached once for validation, while the others are used for training. In total, n(n1)n(n-1) train–validation–test sessions were conducted.
3.3.2 Comparison with State-of-the-Art (RQ3)
This evaluation considered our skeleton-based approach with the ASL-2nd landmark selection strategy, which delivered the highest F1-score on both MINDS-Libras and LIBRAS-UFOP (last rows in Table 1). We then compared our results with Alves et al., (2024) – which inspired this work—, De Castro et al., (2023), Cerna et al., (2021), and the recent work of Naz et al., (2025). We also included a variation of the method by Alves et al., (2024), trained for 30 epochs instead of the original 20, ensuring fairer comparability with our proposal.
3.3.3 Speed-up Evaluation (RQ4)
This experiment compares the time performance of our approach – adopting the ASL-2nd landmark selection strategy – with Alves et al., (2024). While their method relies on OpenPose (Cao et al., , 2021) to extract the signer’s landmarks, our approach employs MediaPipe (Lugaresi et al., , 2019) for faster landmark extraction. For this evaluation, the video “02AlunoSinalizador08-5.mp4” from the MINDS-Libras dataset was selected, as its length (139 frames) is representative of the dataset’s average of 139.41 frames. Landmark extraction, the most time-consuming procedure, was measured for both approaches. Model inference was measured only for Alves et al., (2024), as the time difference was negligible.
3.4 Experimental Platform
The main experiment and comparison with state-of-the-art were run on a native Linux workstation (Ubuntu 22.04). The system features an Intel Core i9-10900KF at 3.70 GHz, 32 GB RAM, and an NVIDIA GeForce RTX 3060 GPU with 12 GB of VRAM. The software stack comprised Python 3.10.12 and PyTorch 2.0.1 with CUDA 12.1. The speed-up evaluation used an Intel Core i7-13700K at 5.4 GHz, with 32 GB of DDR4 RAM at 3200 MT/s, and an NVIDIA GeForce RTX 4070 GPU. Source code, landmarks data, and pretrained checkpoints will be released upon acceptance of this work.
Table 1: Results of the main experiment: comparative view of the landmark subsets. The last column reports the improvement (in percentage points) obtained with the spline-based imputation procedure. Values are presented as averages followed, in parentheses (when applicable), by the standard deviation.
Dataset Land. subset Accuracy Precision Recall F1-score F1-score imp. (p.p.)
MINDS-Libras All 0.70 (0.07) 0.71 (0.06) 0.69 (0.07) 0.66 (0.07) 6
Laines et al., (2023) 0.95 (0.04) 0.95 (0.05) 0.95 (0.04) 0.94 (0.04) 4
Arcanjo et al., (2024) 0.95 (0.05) 0.96 (0.03) 0.94 (0.04) 0.94 (0.05) 4
ASL-1st 0.95 (0.05) 0.95 (0.05) 0.95 (0.05) 0.94 (0.05) 6
ASL-2nd 0.94 (0.04) 0.95 (0.05) 0.95 (0.04) 0.94 (0.05) 4
LIBRAS-UFOP All 0.72 (0.04) 0.74 (0.02) 0.72 (0.04) 0.69 (0.04) 18
Laines et al., (2023) 0.88 (0.05) 0.90 (0.04) 0.88 (0.05) 0.87 (0.05) 17
Arcanjo et al., (2024) 0.91 (0.04) 0.92 (0.04) 0.91 (0.04) 0.90 (0.04) 5
ASL-1st 0.87 (0.05) 0.89 (0.04) 0.87 (0.05) 0.86 (0.05) 15
ASL-2nd 0.91 (0.04) 0.93 (0.03) 0.91 (0.04) 0.91 (0.04) 6
Values highlighted in bold indicate the highest average metric value for a given a dataset.
4 Results and Discussion
The following sections address, individually, the three experiments conducted in this work.
4.1 Main Experiment
Table 1 presents a comparative view of the recognition performance for the five landmark subsets considered in this work. For both datasets, using all landmarks provided by MediaPipe (All) yielded the worst results. For MINDS-Libras, the other landmark subsets performed very similarly, with small fluctuations in the average metrics and standard deviation values (within $\pm$2 p.p.). Nonetheless, a more pronounced contrast was observed for LIBRAS-UFOP. For instance, ASL-2nd achieved an F1-score of 0.91, which is 5 p.p. higher than ASL-1st and 4 p.p. higher than Laines et al., (2023). In summary, ASL-2nd and Arcanjo et al., (2024) presented the more consistent performance in terms of F1-score.
The reduction in input dimensionality (i.e., fewer landmarks) appears to enhance generalization while preserving the essential articulatory cues for isolated sign recognition. Unlike the baseline (All), the evaluated subsets focus primarily on pose and hand landmarks, with reduced or no reliance on dense facial meshes. This design helps mitigate noise caused by factors such as longer camera distances, illumination variability, and occasional occlusions—known to negatively affect facial landmark quality.
The last column in Table 1 shows the performance improvement (in terms of F1-score) resulting from applying the spline-based imputation procedure. Overall, the imputation enhanced the performance of all subsets by at least 4 p.p. The differences were more pronounced for LIBRAS-UFOP, where improvements of over 15 p.p. were observed for three subsets. The obtained results reinforce the need for post-processing the landmarks, as their detection with MediaPipe can be highly unstable.
Table 2: Results of the comparative evaluation with state-of-the-art methods. Values are presented as averages followed, in parentheses (when applicable), by the standard deviation.
Method Year of pub. Dataset Lopo eval.? Accuracy Precision Recall F1-score
Ours (ASL-2nd) MINDS-Libras 0.94 (0.04) 0.95 (0.05) 0.95 (0.04) 0.94 (0.05)
Alves et al., (2024) (30 epochs) 2024 0.94 (0.04) 0.94 (0.05) 0.94 (0.04) 0.93 (0.04)
Alves et al., (2024) 2024 0.93 (0.05) 0.94 (0.05) 0.93 (0.05) 0.93 (0.05)
Naz et al., (2025) 2025 0.97 (0.01) - - -
De Castro et al., (2023) 2024 0.91 (0.07) - - 0.90 (0.08)
Passos et al., (2021) 2021 0.85 (0.02) - - -
Ours (ASL-2nd) LIBRAS-UFOP 0.91 (0.04) 0.93 (0.03) 0.91 (0.04) 0.91 (0.04)
Alves et al., (2024) (30 epochs) 2024 0.82 (0.03) 0.84 (0.03) 0.82 (0.03) 0.80 (0.04)
Alves et al., (2024) 2024 0.82 (0.04) 0.83 (0.05) 0.81 (0.05) 0.80 (0.05)
Naz et al., (2025) 2025 0.89 (0.04) - - -
De Castro et al., (2023) 2024 0.74 (0.04) - - 0.71 (0.05)
Passos et al., (2021) 2021 0.65 (0.04) - - -
Cerna et al., (2021) 2021 0.74 (0.03) - - -
Values highlighted in bold indicate the highest average metric value for a given a dataset.
The ‘-” marker indicates that the metric was not reported in the original publication.
4.2 Comparison with State-of-the-Art
Table 2 presents the results of the comparative evaluation with the literature. Results for Cerna et al., (2021) are reported only for LIBRAS-UFOP. On MINDS-Libras, our accuracy was comparable to Alves et al., (2024), who slightly benefited from training for 30 epochs instead of the original 20. Although Naz et al., (2025) achieved the highest accuracy, their evaluation relied on a less robust cross-validation protocol rather than LOPO.
On LIBRAS-UFOP, we outperformed the competing methods in all metrics, including Naz et al., (2025) (+2 p.p. in accuracy), which employed a less robust cross-validation protocol. Compared to Alves et al., (2024), our method achieved a substantially higher performance across all metrics (+11 p.p. in F1-score). These results demonstrate that a properly selected subset of landmarks, combined with an imputation procedure for missing landmarks, improves accuracy even when using a lighter framework for landmark extraction. The next section provides an alternative perspective on the gains, focusing on the speed-up achieved by our method.
4.3 Speed-up Evaluation
Table 3: Five runs of the recognition pipeline using a sample video from MINDS-Libras. The columns report the times for our approach and for Alves et al., (2024) (both in terms of landmark extraction), followed by the model inference time. The bottom row shows the mean (standard deviation).
# Ours Alves et al., (2024) Inf. time
1 4.612 25.464 0.946
2 4.399 23.385 0.863
3 4.457 38.358 0.861
4 4.421 25.115 0.896
5 4.297 31.524 0.870
Mean (SD) 4.437 (0.102) 28.769 (5.528) 0.887 (0.032)
Table 3 presents the results of the speed-up evaluation. Notably, the average landmark extraction time in our pipeline is 4.4 seconds, compared to 28.7 seconds for the competing method, representing a speed-up nearly 6.7 times. When including the inference time, which better reflects the overall pipeline execution, the observed speed-up was approx. 5.6 times. This result, in conjunction with the accuracy improvement, reinforces the contribution of this work.
5 Conclusion and Future Work
In this paper, we investigated the feasibility of lightweight body landmark detection based on MediaPipe for the recognition of isolated signs in LIBRAS. Preliminary investigations using the pipeline proposed by Alves et al., (2024) indicated that simply replacing OpenPose with the MediaPipe Holistic model (using all available landmarks), although improving time efficiency, dramatically hindered recognition performance.
[…]
Future work includes extending this investigation to other sign languages and datasets, allowing for a more comprehensive analysis of the proposed methodology. Additionally, we plan to adapt feature selection techniques to the SLR domain for automatically optimized recognition performance. Finally, we envision extending our method to continuous SLR, which involves processing video streams at the sentence level.
ACKNOWLEDGEMENTS
Omitted due to anonymous submission.